# MOSAIC: Modeling Social AI for Content Dissemination and Regulation in Multi-Agent Simulations

### v1

@@ 3.3 Content Moderation Improves Both Fact-Checking and Engagement
While political misinformation does not spread faster than factual news in the simulations, another important observation is that overall engagement remains significantly lower in the absence of fact-checking. Figure 4 demonstrates that, under No Fact-Checking, even factual posts fail to gain substantial engagement.
This suggests that the agents, when left without explicit fact-checking mechanisms, tend to suppress their interactions altogether, rather than engaging indiscriminately with unreliable content. We hypothesize that this behavior emerges due to the inherent sensitivity of LLM-based agents to political misinformation, leading them to disengage rather than risk amplifying uncertain information.
To address this disengagement issue, we introduce various fact-checking mechanisms and observe that they not only suppress the spread of misinformation but also enhance engagement with factual news. As shown in Figure 3, we observe a pattern in how different content moderation approaches influence the engagement balance between factual content and misinformation over time. Third Party fact-checking emerges as the most effective intervention, establishing a substantial cumulative advantage for factual content that reaches approximately 325 units by the final time step, significantly outperforming all other approaches. Notably, even the No Fact Check condition maintains a positive trajectory, indicating that factual content may have some inherent engagement advantage in the simulated environment, but this advantage is dramatically amplified by active moderation strategies. The consistently positive and increasingly divergent trajectories suggest that moderation effects compound over time, with the gaps between approaches widening progressively rather than stabilizing. This indicates that content moderation not only creates immediate benefits but generates cumulative advantages for factual information that strengthen with continued application, with third-party professional fact-checking demonstrating particular potency in creating a healthy information ecosystem. This is somewhat corroborated by some human user studies showing content moderation increases trust in unflagged content Pennycook et al. (2020).
3.4 Fact-Checking Performance: Hybrid Model Achieves the Best Trade-off
While Third-Party fact-checking achieves the best engagement-separation effect, we evaluate fact-checking effectiveness using precision, recall, and F1-score (Table 2). The results indicate that the Hybrid approach provides the best overall fact-checking performance, achieving a precision of 0.625, recall of 0.6 and an F1 Score of 0.612, highest amongst all three approaches we explored.
This suggests that a combination of community-based and third-party verification leads to the most reliable identification of misinformation, balancing recall (coverage of false news) and precision (accuracy of detection). In contrast, Third-Party fact-checking alone suffers from a lower recall rate (15.6%), indicating that while it is selective in marking posts as misinformation, it may not capture all false claims effectively. The Community-Based model, on the other hand, achieves higher recall but lower precision, leading to moderate overall performance.
4 Excursion: What Makes Certain Users/Content More Popular?
Understanding why certain content gains traction in online spaces is essential for modeling engagement dynamics and intervention strategies. In this section, we analyze the diffusion characteristics of posts, particularly through agents’ reactions to different social media content.
Figure 5: Top 50 users with highest engagement.
@@ 4 Excursion: What Makes Certain Users/Content More Popular?
Figure 6: User engagement best power-law fit.
Experimental Setup
We generate a diverse set of content types, including news, opinion pieces, and user-generated posts, each with varying levels of controversy, sentiment, and source credibility. Agents interact with these posts based on their personas, engagement signals, and memory-based decision-making. In this experiment, the agents are no longer just prompted to react to the social feeds, but they also participate in generating new content that circulates through the network. We combine two simulation runs that host a total of 161 agents and 4,249 posts.
4.1 Findings
Our study investigates the underlying drivers of user engagement in a simulated social media environment by analyzing multiple dimensions including user popularity, persona attributes, content topics, and agent reasoning patterns. We begin by defining user popularity as the sum of followers, likes, shares, and comments, and observe in Figure 5 that user engagement follows a power-law distribution. The best-fitting approximation, $f(x)=120x^{-0.6}$, with an $R^{2}=0.84$ (Figure 6), confirms an adherence to power-law behavior where a small number of users generate most of the engagement. Though our exponent $\alpha=0.60$ is lower than the typical range (1.5–2.5) reported in real-world networks (Muchnik et al., 2013; Bild et al., 2015), it still highlights the skewed nature of content popularity.
We then explore whether user engagement correlates with persona attributes by comparing the top and bottom 50 users across several traits. The Chi-square analysis summarized in Table 3 reveals no statistically significant differences across categories such as age, gender, activity type, hobby, ethnicity, income level, political affiliation, or primary goal. Although medium effect sizes for ethnicity and hobby (Cramer’s V of 0.319 and 0.302, respectively) are noted Cramér (1946), the absence of statistical significance underscores the lack of strong influence of these traits on engagement. Furthermore, the simulation did not include public figures or celebrity-like personas, indicating that even with randomized initialization, a subset of users naturally attracts more attention, emphasizing the unpredictability of engagement in such networks.
To assess whether content topics drive engagement, we applied BERTopic Grootendorst (2022) using all-MiniLM-L6-v2 embeddings Reimers and Gurevych (2019) to cluster post content and computed engagement statistics for each topic. An ANOVA test returned an F-statistic of 0.614 with a p-value of 0.84, indicating no statistically significant relationship between content topics and engagement. This suggests that the semantic content alone does not reliably influence user reactions.
Given the lack of correlation between engagement and demographic or topical properties, we turn to feed prioritization mechanisms. Our system emphasizes recency and follow relationships rather than engagement-based ranking, yet this setup inadvertently creates a feedback loop. Once a user is followed, their content gains more exposure, which can further increase engagement and visibility, reinforcing the user’s popularity. These findings lead us to speculate that the power-law influence distribution may arise from the agents’ behavioral dynamics, such as copying the actions of prior agents, creating preferential attachment loops that reinforce popularity instead of from their profile attributes or the specific content they generate.
@@ 4.1 Findings
Finally, we examine agents’ reasoning traces to better understand engagement behaviors. Table 4 highlights distinct sentiment and motivation patterns associated with different actions. Positive sentiment dominates actions like following (99%), commenting (97%), liking (92%), and sharing (92%), while negative sentiment is prevalent in flagging (71%) and unfollowing (40%). Motivational reasoning varies by action type: flagging stems from quality assessments (49%) and misinformation concerns (22%), while sharing reflects agreement (46%). Likes and comments are driven by social connection and agreement, and following is based on long-term interest. Vocabulary analysis further reveals that engagement types rely on domain-specific lexicons. Yet, there remains a notable disconnect: only 21% of posts have sentiment alignment with agent reasoning, implying that agents’ verbal justifications do not fully capture the deeper influences behind their choices. Despite articulating rationales involving value alignment and relevance, these factors do not reliably predict popularity, further supporting the role of individual context and network dynamics.
Together, these findings paint a complex picture: while agents provide structured reasoning for engagement actions, overall popularity and content virality arise less from user or content attributes and more from emergent social dynamics and structural effects in feed exposure. We provide more extended discussion and analysis in Appendix H.
@@ 5 Background
Behavioral Economics and Persuasion Games.
We model a sequential persuasion game using LLM-powered agents conditioned on fine-grained personas Kamenica and Gentzkow (2011); Gentzkow and Kamenica (2017); Acemoglu et al. (2023). These agents operate within a directed social graph and evolve based on memory and social context, enabling the study of online behavior, intervention strategies, and moderation effects.
@@
LLMs have transformed agent-based modeling by enabling context-aware, generative behaviors. While early simulations—such as Schelling’s segregation model Schelling (1971), Sugarscape Epstein and Axtell (1996), and NetLogo-based environments Wilensky (1999)—relied on static heuristics, recent systems like Smallville Park et al. (2023), AgentVerse Chen et al. (2024a), and Chirper Minos (2023) showcase agents with lifelike interactions and social dynamics. However, LLM-driven agents still face challenges like inconsistency and limited long-term reasoning. Our work addresses this by incorporating structured constraints and iterative feedback to enhance reliability for social science research.
@@
False information often spreads more rapidly than truth due to emotional appeal and engagement-driven algorithms Vosoughi et al. (2018); Pennycook and Rand (2021); Solovev and Pröllochs (2022). Existing responses—third-party fact-checking Raghunath and Malik (2024); Patel (2024), algorithmic detection, and crowdsourced moderation like Community Notes [^6] https://communitynotes.x.com/guide/en/—each face limitations in scalability, accuracy, or bias Zannettou et al. (2019); Panizza et al. (2023). We use LLM-based simulations to evaluate these approaches in controlled settings, comparing their effectiveness and exploring hybrid strategies.
@@
Simulations have long supported decision-making in fields like epidemiology and public policy Currie et al. (2020); Axtell and Farmer (2022); Qu and Wang (2024). In the context of social media governance, LLM-driven simulations offer a novel testbed for assessing content moderation and algorithmic interventions before deployment Charalabidis et al. (2011); Landau et al. (2024). Our framework enables scalable experimentation with regulatory strategies, contributing to ongoing efforts in algorithmic auditing and platform accountability.
@@ 6 Conclusion
Our study introduces a novel multi-agent generative AI simulation to model content diffusion, engagement, and misinformation dynamics in social networks. Using LLM-driven agents, we bridge social science and computational modeling with high realism. We find that hybrid content moderation—combining community-based and third-party fact-checking—best balances misinformation reduction and user engagement. Notably, LLM agents tend to avoid unverified content, likely due to safety training, and misinformation did not spread faster than factual news, unlike in human studies. Engagement followed a power-law distribution, with few users driving most activity. However, user attributes and content topics were weak predictors, highlighting the complexity of online ecosystems. Agent reasoning showed a gap between stated motivations and actual behavior, suggesting network effects shape engagement more than demographics or content.
@@
Our findings are subject to several limitations, particularly in the scale of our experiments. First, the limited number of human participants, especially from minority demographic groups, restricts the statistical power of our conclusions. Expanding participant diversity would enable a more robust analysis of how alignment between real and simulated social interaction patterns varies across demographics. Second, our content moderation experiments were conducted at a relatively small scale, which may have constrained the emergence of complex behaviors. Running these experiments at a larger scale could uncover additional dynamics not captured in the present study. Third, the fact-checking agents in our simulations lacked access to live web search, limiting their ability to verify claims against up-to-date or external information sources. Enhancing their access to real-time data could significantly improve their reliability and utility. Finally, we observe a gap between agents’ explicit explanations for their actions and the collective reaction patterns that emerge in the system. The root causes of this misalignment remain unclear and warrant further investigation, potentially involving a deeper analysis of agent modeling assumptions or social influence mechanisms.
@@ Acknowledgment
This work was supported in part by research credits generously provided by OpenAI. In addition, the authors would like to thank Professor Jacob Andreas for his valuable feedback and insightful suggestions.
Appendix A Persona Generation Details
A.1 Persona Replication from Human Survey
A.2 Synthetic Persona from Agent Bank
A.3 User Generation and Instantiation
Persona Generation
As illustrated in Figure 1, personas are generated using a combination of predefined questions and sampling from probabilistic distributions stored in an agent_bank profile collection, inspired by Park et al. (2024). Key demographic attributes such as age, gender, ethnicity, and primary language are assigned probabilistically to mirror real-world distributions. For instance, age follows a normal distribution centered around 35 years, while other attributes are sampled based on predefined probabilities. We disclose these questions and describe more detail of the methodology in Appendix A. After synthesizing the structured profiles for agents, we construct a natural language description for each of them. This process leverages a mixture of deterministic rules and LLM-based augmentation using GPT-4o Hurst et al. (2024) to enhance diversity and realism.
User Instantiation
Once personas are generated, they are instantiated as agent users within the simulation. Each agent is assigned a unique user ID and a persona profile that includes background details and interest labels. The relational database serves as the backbone for recording agent activities, ensuring persistent storage of interactions, post engagements, and behavioral updates. This database facilitates dynamic user tracking and enables post-simulation analysis of engagement trends and content spread. We provide more implementation details in Appendix D.
Appendix B Details of the Human Study
The study was open to 20,240 eligible participants from a larger Prolific population of 232,330, and we collected 204 valid responses from eligible participants. The survey was conducted via Prolific to collect responses from U.S.-based participants fluent in English. Participants were asked to complete a 12-minute survey assessing their demographic characteristics and social media interactions. The survey, hosted on Google Forms, required no software downloads or special device features and was accessible via mobile, tablet, or desktop. Participant recruitment applied custom screening for language, political spectrum, vaccine opinion, and prior participation, ensuring a targeted sample. Responses were collected using Prolific ID via a question at the start of the form, and participants received a completion code upon finishing. Compensation was set at \$2.40 per participant, equivalent to \$12.00/hour, and submissions were manually reviewed before approval. The median completion time was approximately 14.5 minutes. All members on our research team have obtained IRB approval before the human study was conducted. Our study costs a total of \$480 for participant payment and \$160 of platform fee. Figure 7 shows the complete breakdown of the 9 key demographic distributions of the 204 human participants.
Figure 7: Demographic Distributions of Study Participants.
B.1 Per-Demographic Attribute Engagement Pattern
We also analyzed the reaction patterns between human participants and persona-driven agents grouped by specific demographic attributes such as age, gender, income, ethnicity, etc, as shown in Table 1.
Overall, out of 52 demographic groups analyzed, 14 show significant differences in one or more engagement types, while 38 do not. The criteria for significance were based on a p-value of less than 0.05 in statistical comparisons. The results suggest that agents may be more adept at simulating the engagement patterns of "common" or more broadly represented demographic groups in LLM pretraining data, as indicated by the lack of significant differences in many of these groups.
Appendix C An Extended Version of Related Work
Behavioral Economics and Persuasian Games.
AI-Driven Social Simulations and Generative Agents.
Misinformation Spread and Fact-Checking Mechanisms.
Simulations as Tools for Policy and Platform Governance
Appendix D Database Schema
Appendix E Agent Action Space
Figure 8: Agent’s action space.
E.1 Agent Decision-Making Process
Feed Presentation
Memory and Reflection Module
We implement an AgentMemory module which manages the memory and reflection capabilities of each agent. Memories are categorized into interactions (e.g., past engagements) and reflections (high-level insights derived from past behaviors). Each piece of memory is assigned an importance score, which decays over time unless reinforced by further interactions. The decay function ensures that long-term behaviors emerge naturally based on experience. Please refer to Appendix G for more details of the Memory module.
Agent Decision-Making and Action Execution
Appendix F Detailed Experiment Configuration
Appendix G Details of the Memory Module
Memory relevance is computed as:
[…]
The decay factor is defined as:
$\text{Decay}=\max(0,\text{PrevDecay}-\alpha\Delta t)$
Importance Scoring.
  • Emotional: love, hate, angry, happy, sad
  • Action: achieved, failed, learned, discovered
  • Relationship: friend, follow, connect, share
  • Goal: objective, target, aim, purpose
Let $k$ be the number of keyword matches in the memory content. Then:
$\text{Importance}=\min(1.0,0.5+0.1k)$
Appendix H Extended Discussion on Content Popularity
H.1 Power-Law Distribution of User Popularity

### v2

@@ 3.3 Content Moderation Improves Both Fact-Checking and Engagement
While political misinformation does not spread faster than factual news in our simulations, overall engagement remains low without any fact-checking. Fig. 4 shows that, in the No Fact-Checking baseline condition, agents interact very little with both true and false posts. We hypothesize that LLM-driven agents, when uncertain about a post’s veracity, choose to disengage rather than risk amplifying misleading content.
Surprisingly, with the community-based, third-party, and hybrid fact-checking, we find that each not only suppresses misinformation (Tab. 1) but also increases engagement with factual content. In particular, Third-Party verification produces the largest cumulative advantage for real news, achieving $\sim$325 interactions by the final time step (Fig. 3). Even the unmoderated baseline shows a modest upward trend for factual posts, hinting at an inherent bias toward accurate information in our network.
The steadily widening gaps between the third-party/hybrid approaches indicate that certain content moderation generates compounding benefits for factual engagement over time. Among the simulated strategies, fact-checking by third parties proves most effective at building a healthy information ecosystem—an effect that aligns with human studies showing increased trust in unflagged content Pennycook et al. (2020). All curves in Fig. 4 represent metrics averaged across our three LLM agents (DeepSeek-V3, GPT-4o, and Claude-3.7-Sonnet). Interaction counts including the number of shares, likes, comments, and notes, are aggregated at each time step to highlight overall platform dynamics under each moderation strategy. This high-level view illustrates consistent patterns in how content moderation shapes engagement. For a detailed, model-by-model breakdown and finer-grained metric analysis, please see Appendix J.
3.4 Fact-Checking Performance: Hybrid Shows Superior Balance Across Models
We evaluate fact-checking effectiveness using precision, recall, and F1 score (See Tab. 1) across all three models—DeepSeek-V3 Liu et al. (2024), GPT-4o Hurst et al. (2024), and Claude-3.7-Sonnet Anthropic (2025). We find that LLM-based fact-checkers consistently perform best in the Hybrid Fact-Checking setting. Claude outperforms GPT-4o and DeepSeek, achieving a F1 score of 0.895, with perfect precision (1.0) and high recall (0.810).
Third Party Checking Ablation
For GPT-4o, we conduct an ablation over different third-party fact-checking configurations. Increasing the number of posts reviewed per step and enabling web search both lead to notable improvements in recall and overall F1 score. The 5-post offline setting yields a modest F1 score of 32.6%, with low precision (21.9%) and moderate recall (63.6%). Switching to the 5-post online setup boosts recall to 100%, indicating that the system catches all misinformation posts in this condition, though precision remains low (22.2%), resulting in a slightly improved F1 score of 36.4%. The best third-party performance is observed in the 10-post online setup, which balances both precision (51.4%) and recall (90.0%), achieving an F1 score of 65.5%. These results suggest that increasing review coverage and leveraging web search significantly enhance the effectiveness of third-party fact-checking.
3.5 What Drives Users/Content Popularity?
@@
Figure 5: User engagement’s best power-law fit. Engagement is defined as the sum of reposts, likes, and comments received by the user.
Understanding what makes some content more popular is key for modeling engagement and intervention. In this standalone study, we simulate 161 agents over 4,249 posts of news and user-generated content, and let agents both react to and generate content based on personas, engagement cues, and memory-based decisions.
@@
Findings
Engagement follows a heavy-tailed power law ($R^{2}=0.84$, $\alpha=0.60$; Figure 5), even when user persona and the network graph are initialized randomly. Few users attract most interactions despite a lower exponent than real networks (Muchnik et al., 2013; Bild et al., 2015). Most persona demographic attributes show no significant effect on popularity (Tab. 2), even though out of 8 attributes, ethnicity and hobby have medium effect sizes (Cramer’s $V=0.319,0.302$) Cramér (1946). Topic clusters via BERTopic on all-MiniLM-L6-v2 Grootendorst (2022); Reimers and Gurevych (2019) embeddings reveal no engagement differences (ANOVA $F=0.614$, $p=0.84$).
Finally, we examine agents’ own reasoning traces to better understand engagement behaviors. Interestingly, agent reasoning traces show that sentiment and motivations do not always align with action types (e.g., positive sentiment in follows, likes, shares; quality or misinformation concerns in flags). As shown in Tab. 3, 13.5% of “flags” are coded as positive even though flag reasons are exclusively negative (information value, misinformation). The verbal reasoning for the “unfollow” actions is 20% positive though it is intuitively a negative reaction towards other users. “ignore” omits any neutral reasoning and a 77.8% positive rate in reasoning, which also makes their verbal explanations inconsistent with the actual action. Following these inconclusive analyses, we speculate that our recency-and-follow-based feed ranking creates preferential attachment: once followed, a user’s posts gain visibility, reinforcing their popularity. We provide further analyses in Appendix A.
@@ 4 Background
LLMs have transformed agent-based modeling by enabling context-aware, generative behaviors. Early simulations such as Schelling’s segregation model Schelling (1971), Sugarscape Epstein and Axtell (1996), and NetLogo-based environments Wilensky (1999) rely on static heuristics; recent systems Park et al. (2023); Chen et al. (2024a); Piao et al. (2025); Minos (2023) showcase agents with lifelike interactions and social dynamics. However, LLM-driven agents still face challenges like inconsistency and limited long-term reasoning. Our work addresses this by incorporating structured constraints and iterative feedback to enhance reliability for social science research.
@@
False information often spreads more rapidly than truth due to emotional appeal and engagement-driven algorithms Vosoughi et al. (2018); Pennycook and Rand (2021); Solovev and Pröllochs (2022). Existing responses—third-party fact-checking Raghunath and Malik (2024); Patel (2024), algorithmic detection, and crowdsourced moderation like Community Notes [^10] https://communitynotes.x.com/guide/en/—each face limitations in scalability, accuracy, or bias Zannettou et al. (2019); Panizza et al. (2023). We use simulations to evaluate these approaches in controlled settings, comparing their effectiveness and exploring hybrid strategies.
@@
Simulations have long supported decision-making in fields like epidemiology and public policy Currie et al. (2020); Axtell and Farmer (2022); Qu and Wang (2024); Cai et al. (2025). In the context of social media governance, LLM-driven simulations offer a novel testbed for assessing content moderation and algorithmic interventions before deployment Charalabidis et al. (2011); Landau et al. (2024); Qiu et al. (2025); Gu et al. (2025). Our framework enables scalable experimentation with regulatory strategies, contributing to ongoing efforts in algorithmic auditing and platform accountability.
@@ 5 Conclusion
Our study introduces a novel generative multi-agent simulation to model content diffusion, engagement, and misinformation dynamics in social networks. Our proposed moderation, combining community-based and independent fact-checking, balances misinformation reduction and user engagement. Notably, LLM agents tend to avoid unverified content, likely due to safety training, and misinformation did not spread faster than factual news, unlike in human studies. Engagement followed a power-law distribution, with few users driving most activity. However, user attributes and content topics were weak predictors, highlighting the complexity of online ecosystems. Agent reasoning showed a gap between stated motivations and actual behavior, suggesting future work on the faithfulness of agentic reasoning.
@@
Our findings are subject to several limitations, particularly in the scale of our experiments. First, the limited number of human participants, especially from minority demographic groups, restricts the statistical power of our conclusions. Expanding participant diversity would enable a more robust analysis of how alignment between real and simulated social interaction patterns varies across demographics. Second, our content moderation experiments were conducted at a relatively small scale, which may have constrained the emergence of complex behaviors. Running these experiments at a larger scale could uncover additional dynamics not captured in the present study. Finally, we observe a gap between agents’ explicit explanations for their actions and the collective reaction patterns that emerge in the system. The root causes of this misalignment remain unclear and warrant further investigation, potentially involving a deeper analysis of agent modeling assumptions or social influence mechanisms.
@@ Acknowledgment
This work was supported in part by research credits generously provided by OpenAI. In addition, the authors would like to thank Jacob Andreas for his valuable feedback and insightful suggestions.
Appendix A Extended Discussion on Content Popularity
A.1 Network Properties: Centrality and Clustering Metrics
We analyze the structural role of each agent in the directed follow-network $G=(V,E)$ with $N=|V|$ nodes by computing degree, betweenness, closeness, and eigenvector centralities, as well as the local clustering coefficient and global transitivity.
Degree centrality quantifies how many direct connections a node has relative to the maximum possible. For node $v$ with total degree $\deg(v)=k_{v}$ (in- plus out-degree), we define
$C_{D}(v)=\frac{k_{v}}{N-1},$
and compute the average degree centrality as
$\overline{C_{D}}=\frac{1}{N}\sum_{v\in V}C_{D}(v).$
Betweenness centrality measures the fraction of shortest directed paths between all ordered pairs $(s,t)$ that pass through $v$. Let $\sigma_{st}$ be the number of shortest paths from $s$ to $t$, and $\sigma_{st}(v)$ those that pass through $v$. Then
$C_{B}(v)=\sum_{\begin{subarray}{c}s,t\in V\\ s\neq t\neq v\end{subarray}}\frac{\sigma_{st}(v)}{\sigma_{st}},$
and its normalized average is
$\overline{C_{B}}=\frac{1}{N\,(N-1)\,(N-2)}\sum_{v\in V}C_{B}(v).$
Closeness centrality reflects how near a node is to all others based on shortest-path distances $d(v,u)$. We set
$C_{C}(v)=\frac{N-1}{\sum_{u\in V\setminus\{v\}}d(v,u)},$
and average as
$\overline{C_{C}}=\frac{1}{N}\sum_{v\in V}C_{C}(v).$
Eigenvector centrality assigns importance proportional to the centrality of a node’s neighbors. If $\mathbf{A}$ is the adjacency matrix of $G$, we solve
$\mathbf{A}\,\mathbf{x}=\lambda_{\max}\,\mathbf{x},$
and take $C_{E}(v)=x_{v}$, with average
$\overline{C_{E}}=\frac{1}{N}\sum_{v\in V}x_{v}.$
The local clustering coefficient of node $v$ measures the density of edges among its $k_{v}$ neighbors. If $e_{v}$ is the number of edges between neighbors, then
$C(v)=\frac{2\,e_{v}}{k_{v}(k_{v}-1)},$
and the network-wide clustering is
$\overline{C}=\frac{1}{N}\sum_{v\in V}C(v).$
Finally, transitivity (global clustering) is the ratio of closed triplets (triangles) to all connected triplets:
$T=\frac{\sum_{v\in V}2\,e_{v}}{\sum_{v\in V}k_{v}(k_{v}-1)}.$
In these definitions, $d(v,u)$ denotes the shortest directed-path length from $v$ to $u$, $\sigma_{st}$ the total number of shortest paths from $s$ to $t$, and $\sigma_{st}(v)$ those passing through $v$. All sums run over $V$ unless noted otherwise.
In our simulated community of 161 agents engaging with 4,249 pieces of content, the resulting follow-network comprises 358 directed ties, indicating a modest level of connectivity driven by personas and memory-based decisions. An average degree centrality of 0.1105 shows that each agent follows roughly 11% of the population, while the low average betweenness centrality (0.0128) suggests that few agents act as indispensable bridges. With an average closeness centrality of 0.1367, most agents remain just a few steps apart, supporting rapid information flow, and the modest eigenvector centrality (0.0433) reveals that influence is fairly distributed rather than monopolized by a handful of hubs. Finally, a clustering coefficient of 0.205 and network transitivity of 0.189 point to moderate local grouping without excessive fragmentation—together painting a picture of a network that balances cohesion and reach in propagating popular content.
A.2 Power-Law Distribution of User Popularity